OpenStudy (anonymous):

can someone help me factor x^3-25x?

OpenStudy (anonymous):

the GCF = x so: = x(x^2 - 25) = x(x+5)(x-5)

OpenStudy (anonymous):

x(x+5)(x-5)

OpenStudy (anonymous):

the bracket contains difference of two squares a^2-b^2 = (a-b)(a+b)

OpenStudy (anonymous):

(x-5) x (x+5)